OpenStudy (blackops2luvr):

What is the value of the expression (–2)^5? A. 10 B. 32 C. –10 D. –32

OpenStudy (anonymous):

D

OpenStudy (blackops2luvr):

@study-mode

OpenStudy (ajprincess):

(-2)^5=(-2)*(-2)*(-2)*(-2)*(-2)=? (-)*(-)=(+) (-)*(+)=(-)

OpenStudy (anonymous):

D -32

OpenStudy (blackops2luvr):

What is the value of the expression –3x^3 when x = –3? A. –81 B. –27 C. 27 D. 81

OpenStudy (blackops2luvr):

@phi

OpenStudy (blackops2luvr):

everyone follow me